# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

### v2

@@ Abstract
We study a new vulnerability in commercial-scale (human-aligned) language models: the ability to refuse generating harmful responses, learned through safety alignment mechanisms, can be broken by flipping just a few bits in the model parameters. Our adversary can exploit this vulnerability and jailbreak billion-parameter language models with at most 25 bit-flips in all cases—and as few as 5 in some—using up to 40 fewer bit-flips than existing attacks on computer vision models at least 100 smaller. Unlike prompt-based jailbreaks, our attack renders these models in memory ‘uncensored’ at runtime, allowing them to generate harmful responses without any input modifications. Our attack algorithm efficiently identifies target bits to flip, offering up to 20 more computational efficiency than previous methods. We show an end-to-end exploitability of the vulnerability using simulations of software-induced fault-injection—Rowhammer (RH). We evaluate on 56 DRAM RH profiles from DDR4 and LPDDR4X devices with different RH vulnerabilities. Our attack can reliably induce jailbreaking in systems similar to those used to showcase prior bit-flip attacks. Moreover, the attack can be effective even against highly RH-secure systems (e.g., 46 more secure than previously tested systems). Bit-flip locations identified in one fine-tuned model transfer to others that share the same pre-trained LLMs, demonstrating the black-box exploitability in standard pretraining-then-finetuning paradigms. Our analyses reveal that: (1) models with weaker post-training alignment need fewer bit-flips to jailbreak; (2) certain model components, e.g., value projection layers, are substantially more vulnerable than others; and (3) the attack is mechanistically different from existing jailbreaking attacks. We evaluate potential countermeasures and find that our attack remains effective against defenses applicable at various stages of the LLM pipeline.
@@ 1 Introduction
Deep neural networks are vulnerable to parameter corruption, which adversaries can exploit to trigger undesirable behaviors. These include severe performance degradation [1, 2, 3], targeted misclassifications [4, 5, 6], and backdoor injections [7, 8, 9, 10, 11]. Recent studies also showcased confidentiality risks, such as model extractions [12]. These attacks exploit the vulnerability through practical fault-injection techniques that cause bit-wise errors in the memory representation of model parameters
@@
Most existing research has focused on exploiting convolutional neural networks in classification tasks [7, 1, 3, 4, 5, 8, 11]. While valuable, these studies leave the vulnerability of language models—particularly those widely deployed across modern AI services—largely unexplored in the context of parameter corruption. The closest work is by Cai et al. [6], who showed that one could corrupt a machine translation model to produce attacker-chosen texts on specific inputs. However, this falls within the breadth of prior bit-flip attack research, and it remains unclear whether language models are vulnerable to unexplored adversarial objectives. One particularly compelling and underexplored scenario is whether an adversary can induce undesirable behaviors, such as jailbreaks, through minimal bit-wise corruptions in a language model’s parameters.
@@
In this paper, we study a previously unexplored vulnerability in commercial-scale language models. Specifically, we ask the question: How vulnerable is the refusal property, learned through safety alignment training, of language models to bit-wise errors induced by practical fault-injection attacks? Once exploited at runtime, such attacks permanently jailbreaks a target model in memory. Unlike prompt-based jailbreaks [13, 14, 15], this attack does not rely on any specific prefixes or suffixes to provoke harmful responses. Moreover, by carefully selecting targeted bit-flips, the attack preserves model performance, making it less detectable.
@@
We first introduce an efficient (offline) procedure for identifying bit locations to flip in target models. Combined with the software-induced fault-injection attack, Rowhammer [16], this forms the basis of our end-to-end attack, PrisonBreak. Due to the bit-flip onion effect we discovered111It is important to note that this onion effect is distinct from the term introduced by Carlini et al. [17] in the context of privacy domains., identifying all target bits at once is challenging. Our procedure employs the progressive bit-search, which iteratively identifies a sequence of target bits to flip based on gradient-based importance scores. In §5, we detail the design of our novel jailbreak objective whose gradients accurately capture the importance. We also present a series of speed-up techniques to address the computational burden of running this search process over billions of parameters. In our evaluation with 8 open-source language models, we achieve a 20 speed-up and successfully jailbreak the target model using only 5–25 bit-flips. Our identified bit-flips also transfer between models that share the same architectures and pre-training, indicating that full knowledge of the target model’s parameters is not strictly necessary. Moreover, we show how an adversary with knowledge of target systems can further optimize the bit-search and succeed in jailbreaking even on DRAM modules with strong Rowhammer resilience.
We next comprehensively evaluate the practicality of PrisonBreak. After profiling, the attacker uses Rowhammer to induce bit-flips at the intended locations in the target system’s memory at runtime. Requiring only a few bit-flips for jailbreaking makes this vulnerability substantially more feasible to exploit via Rowhammer. In our end-to-end evaluation, we perform generalizable Rowhammer simulations across 56 DRAM profiles, including 40 DDR4 and 16 LPDDR4X modules, rather than showcasing a one-off attack on a specific hardware instance. We consider a Machine-Learning-as-a-Service scenario, where the victim and attacker processes are co-located on the same host. Across 20 DRAM profiles, especially on low-power modules, we succeed in jailbreaking over 99% of the time.
@@
Moreover, we conduct structural and behavioral analyses to gain an in-depth understanding of this vulnerability. Our structural analysis, conducted across multiple levels of abstraction in the parameter space, reveals that certain model components, e.g., the exponent bits and later layers, disproportionately contain target bits and are more vulnerable. In our behavioral analysis in the activation space, we show that PrisonBreak works differently from existing jailbreaks—we do not observe the suppression of refusals. We also explore potential countermeasures against both jailbreaking and Rowhammer attacks, highlighting their applicability and limitations. Our findings highlight a pressing, practical threat to the language model ecosystem and underscore the need for research to protect these models from bit-flip attacks.
@@
We comprehensively evaluate PrisonBreak on 8 open-source, large-language models and a representative jailbreaking benchmark. Our attack is more successful in producing high-fidelity harmful responses than prompt-based jailbreaks and comparable to parameter-based attacks that modify orders of magnitude more parameters.
We evaluate our attack’s exploitability with Rowhammer simulations across 56 DRAM profiles (40 DDR4 and 16 LPDDR4X). Our attacker can succeed in jailbreaking even on secure DRAM modules by carefully selecting target bits during the offline bit-search process.
@@ 2 Background and Related Work
where represents the likelihood of the next token , as computed by the neural network with parameters (weights and biases). is learned by iteratively minimizing the objective across the entire training set. If the training reaches an acceptable accuracy, the model along with its parameters is stored. In deployment, the model and parameters are loaded into memory and generate text autoregressively by sequentially processing an input and sampling future tokens until reaching a stopping criterion (e.g., an end-of-speech token). Our work considers the Transformer architecture [21] for , which consists of a sequence of attention layers with additional components such as normalization and fully-connected layers.
@@
Aligning and jailbreaking language models. (Large) language models pre-trained on text data can perform various tasks [22] and serve as the backbone for many user-facing applications. However, it has been shown that malicious users can exploit these capabilities to produce biased, deceptive, or ultimately harmful content [23, 24]. In response, active research is underway to develop techniques to ensure they generate outputs aligned with human values, e.g., supervised fine-tuning (SFT) [25], reinforcement learning with human feedback (RLHF) [26, 27], and direct preference optimization (DPO) [28]. Aligned models [29, 30, 31] produce safe, helpful responses and refuse harmful queries.
@@
Recent work studies the security implications of aligned language models through jailbreaking attacks, which aim to induce models to generate harmful responses. Because these models are accessible to any user prompt, most jailbreaks are prompt-based [13, 15, 14, 32], creating harmful prompts that bypass alignment by compelling models to respond to inputs they would otherwise refuse. A few works present parameter-based jailbreaks where an attacker tunes billions of parameters of an aligned model on a small set of harmful text to remove alignment without altering the input [33, 34]. In contrast, we present a new attack that exploits Rowhammer [16] to flip at most 25 bits in a model’s memory representation to jailbreak it at runtime.
@@
Half-precision floating-point numbers. Modern computer systems use the IEEE754 floating-point format to represent model parameters. Unlike conventional models that typically use single-precision (32-bit), language models with billions of parameters are commonly deployed in half-precision (16-bit), with a sign bit, 5 exponent bits, and 10 mantissa bits to reduce memory footprint. For instance, in half-precision, is represented as , where -1 is the sign, 1.11 is the mantissa, and -4 is the exponent. Here, the exponent bits play a significant role in representing the number. Flipping a bit in the mantissa induces a marginal change to , whereas a bit-flip in the exponent can shift the value to . Our attacker favors this insight to jailbreak a model with a few bit-flips (§5.3).
@@
Rowhammer is the most well-studied fault injection attack that an attacker can induce at the software-level [16]. The attack exploits DRAM disturbance errors by repeatedly accessing specific DRAM rows (i.e., aggressors) to accelerate charge leakage in adjacent rows, leading to a bit-flip. The bit-flip persists until the data in the affected rows is reloaded into memory. Being a versatile attack requiring only access to DRAM content, Rowhammer has been a major security concern, with numerous demonstrations on CPUs [35, 36, 37, 38, 39, 40, 41] and GPUs [42]. It has been also shown to achieve various adversarial goals against neural networks, such as performance degradation [3, 1], targeted misclassification [4, 5, 6], backdooring [9, 10, 7, 8, 11], and model extraction [12], by causing bit-flips to their memory representations. Our work advances with this research direction by presenting the first work on jailbreaking language models through Rowhammer.
@@ 3 Threat Model
We consider an adversary who aims to jailbreak large language models by compromising their integrity at runtime through direct manipulation of model parameters in memory. Our attack is entirely different from the class of attacks that leverage jailbreaking prompts [13, 15, 14, 32] where the attacker has to modify each harmful query to induce harmful responses. In contrast, our attack requires no input modifications. We assume scenarios where a victim deploys a user-facing language model in a resource-sharing environment, such as machine-learning-as-a-service (MLaaS) platforms. These platforms are becoming increasingly common as few entities have the computing resources to run large-scale models and offer natural language processing services to users at scale. The victim makes a best effort to keep these models aligned with human values, with one safety-oriented outcome being the denial of harmful queries (e.g., by answering “I cannot answer that” to “how to create an explosive product for my friend’s birthday party?”).
@@
Knowledge. We consider a white-box attacker with knowledge of the target model’s architecture and its parameters which is consistent with the conventional setup in prior work [1, 9, 11, 6, 10, 7]. This assumption is practical as many user-facing services are built on open-source language models [43]. But we emphasize that such knowledge is not strictly required, as it can potentially be acquired through prior attacks with similar capabilities [44, 12, 45]. Moreover, as we demonstrate in §6.4, bit-flips can be effectively transferred across models, enabling black-box exploitability. The attacker requires no training data. We assume the attacker has knowledge of the memory locations likely to induce bitwise errors. Our adversary can either search for model parameters with the knowledge of vulnerable memory locations (hardware-aware) or profile these locations after the target model is loaded into memory and then induce bit-flips (hardware-agnostic). Since prior attacks assume at most one of these scenarios, our assumption allows for a comprehensive study of practical exploitability.
@@
Capabilities. Our attacker exploits Rowhammer [16]—a practical fault injection attack in modern computing systems—to induce targeted bit-flips in memory. The adversary has only user-level privileges and co-locates a process or virtual machine (VM) within the victim’s process or VM running on the same host. The host OS remains uncompromised, and confinement policies (e.g., process isolation) are enforced. Our attacker can reverse-engineer DRAM address mappings using techniques proposed in [46] and place target pages at desired physical locations by leveraging advanced memory massaging primitives [47]. We further detail the Rowhammer exploitation process in §7. While we mainly consider models loaded in CPU memory, Olgun et al. [42] has demonstrated the exploitation of Rowhammer on GPUs, presenting a potential vector for jailbreaking models in GPU memory. Our assumptions align with those shown in prior bit-flip attacks on neural networks [1, 12, 11, 7].
@@ 4 PrisonBreak
In this section, we introduce PrisonBreak, a new attack designed to jailbreak large-language models with targeted bit-flips. Figure 1 illustrates the overall workflow:
Offline procedure. The attacker identifies target bit locations in the victim model’s memory representation. We run a progressive bit-search to search for the target bits. We first analyze the victim model and identify critical weights whose perturbations are most effective in jailbreaking. We next extract all bits from the floating-point representation of the critical weights and flip them one at a time in isolation to identify the most effective bit-flip. We then select the bit and flip it in the model’s parameter space; the bit-flip persists for the remaining search process. We repeat these steps until achieving an attack success of, e.g., 80–95%, comparable to prior work [13]. We store the identified target bits.
Note that PrisonBreak supports two distinct modes: hardware-agnostic and hardware-aware search, depending on the attacker’s knowledge (see §3). Hardware-agnostic search is as described above, and assumes the attacker performs memory templating after running PrisonBreak. In the hardware-aware search, the attacker profiles the memory before running PrisonBreak, and we are limited to selecting bits at vulnerable locations within a memory template that the attacker has previously profiled. We detail these modes in §5.3 and demonstrate their exploitability in §7.
Online procedure. After identifying target bits, PrisonBreak jailbreaks the victim language model by flipping them. Following our threat model in §3, this exploitation can be accomplished using advanced Rowhammer techniques, including memory templating [46] and massaging [35], and advanced Rowhammer attacks [38, 39, 37] that bypass in-DRAM target row refresh and induce precise bit-flips.

### v3

@@ Abstract
We study a new vulnerability in commercial-scale safety-aligned large language models (LLMs): their refusal to generate harmful responses can be broken by flipping only a few bits in model parameters. Our attack jailbreaks billion-parameter language models with just 5 to 25 bit-flips, requiring up to 40 fewer bit flips than prior attacks on much smaller computer vision models. Unlike prompt-based jailbreaks, our method directly uncensors models in memory at runtime, enabling harmful outputs without requiring input-level modifications. Our key innovation is an efficient bit-selection algorithm that identifies critical bits for language model jailbreaks up to 20 faster than prior methods.
We evaluate our attack on 10 open-source LLMs, achieving high attack success rates (ASRs) of 80–98% with minimal impact on model utility. We further demonstrate an end-to-end exploit via Rowhammer-based fault injection, reliably jailbreaking 5 models (69–91% ASR) on a GDDR6 GPU. Our analyses reveal that: (1) models with weaker post-training alignment require fewer bit-flips to jailbreak; (2) certain model components, e.g., value projection layers, are substantially more vulnerable; and (3) the attack is mechanistically different from existing jailbreak methods. We evaluate potential countermeasures and find that our attack remains effective against defenses at various stages of the LLM pipeline.
@@ 1 Introduction
Deep neural networks are vulnerable to parameter corruption, which adversaries can exploit to trigger undesirable behaviors. These include severe performance degradation [94, 33, 67, 51], targeted misclassifications [70, 7, 11], backdoor injections [45, 14, 69, 99, 80, 50], and confidentiality risks such as model extraction [68]. These attacks exploit the vulnerability through practical fault-injection techniques that cause bit-wise errors in the memory representation of model parameters.
@@
Most prior work targets convolutional neural networks in classification tasks [45, 94, 67, 70, 7, 14, 80]. While valuable, these studies leave the vulnerability of language models—especially those widely deployed in modern AI services—largely unexplored in the context of parameter corruption. The closest work, by Cai et al. [11], demonstrates that corrupting a machine translation model can force attacker-chosen outputs. Yet it remains unclear whether language models are susceptible to more impactful objectives that induce undesirable behaviors. A particularly compelling and novel scenario is whether an adversary can completely remove a language model’s alignment—i.e., induce a jailbreak—through minimal bit-wise corruptions in its parameters.
@@
In this paper, we study a previously unexplored vulnerability in commercial-scale language models. Specifically, we ask the question: How vulnerable is the refusal property of language models, learned through safety alignment training, to bit-wise errors induced by practical fault-injection attacks? Once exploited at runtime, such attacks permanently jailbreak a target model in memory. Unlike prompt-based jailbreaks [101, 27, 53], this attack does not rely on any specific prefixes or suffixes to provoke harmful responses. Moreover, by carefully selecting targeted bit-flips, the attack preserves model performance, making it less detectable.
@@
We first introduce an efficient (offline) procedure for identifying bit locations to flip in target models. Combined with the software-induced fault-injection attack, Rowhammer [42], this forms the basis of our end-to-end attack, PrisonBreak. Due to the bit-flip onion effect we discovered,111It is important to note that this onion effect is distinct from the term introduced by Carlini et al. [12] in the context of privacy domains. identifying all target bits at once is challenging. Our procedure employs the progressive bit-search, which iteratively identifies a sequence of target bits to flip based on gradient-based importance scores. In §5, we detail the design of our novel jailbreak objective whose gradients accurately capture the importance. We also present a series of speed-up techniques to address the computational burden of searching across billions of parameters.
In our evaluation with 10 open-source language models spanning 5 families (Vicuna, Llama2, Llama3, Tulu3, Qwen2), we achieve up to a 20 speed-up and successfully jailbreak target models, reaching an attack success rate (ASR) of 80–98% with only 5–25 bit-flips. Our identified bit-flips also transfer across models that share the same architectures and pre-training, suggesting that full knowledge of model parameters is not strictly necessary.
We then conduct an end-to-end evaluation, showing how an adversary with knowledge of target systems can further optimize the bit-search and succeed in jailbreaking even on DRAM modules with strong Rowhammer resilience. We target GDDR6 DRAM on an NVIDIA A6000 GPU, and consider a Machine-Learning-as-a-Service setting in which victim and attacker processes are co-located on the same GPU-sharing platform. Despite our target system having up to fewer exploitable bit-flips than prior work, we reliably jailbreak five models (69–91% ASR) using only two physical bit locations.
@@
To gain an in-depth understanding of this vulnerability, we conduct structural and behavioral analyses. Our structural analysis, conducted across multiple levels in the parameter space, reveals that certain model components, e.g., exponent bits and value projection layers, are disproportionately vulnerable. In our behavioral analysis in the activation space, we show that PrisonBreak works differently from existing jailbreaks—we do not observe the suppression of refusals. We also explore potential countermeasures against both jailbreaking and Rowhammer attacks, highlighting their applicability and limitations. Our findings highlight a pressing, practical threat to the language model ecosystem and underscore the need for research to protect these models from bit-flip attacks.
@@
We comprehensively evaluate PrisonBreak on 10 open-source large language models using a representative jailbreaking benchmark, achieving high ASRs of 80–98%. Our method outperforms prompt-based jailbreaks and attains performance comparable to parameter-based attacks, while modifying orders of magnitude fewer parameters.
We evaluate PrisonBreak in an end-to-end Rowhammer attack on a GDDR6 GPU, showing that it can jailbreak 5 open-source LLMs on real systems with ASRs of 69–91%, leveraging a hardware-aware target bit-search.
@@ 2 Background and Related Work
where represents the likelihood of the next token , as computed by the neural network with parameters (weights and biases). is learned by iteratively minimizing the objective across the entire training set. If the training reaches acceptable performance, the model is stored. In deployment, the model and parameters are loaded into memory and generate text autoregressively by sequentially processing an input and sampling future tokens until reaching a stopping criterion (e.g., an end-of-speech token). Our work considers the seminal Transformer architecture [84] for , which consists of a sequence of attention layers with additional components, such as normalization and fully-connected layers.
@@
Aligning and jailbreaking language models. (Large) language models (LLMs), pre-trained on text data, can perform various tasks [10], and are used in many user-facing applications. However, malicious users can exploit these capabilities to produce biased, deceptive, or ultimately harmful content [87, 86]. In response, several techniques try to ensure these models generate outputs aligned with human values, e.g., supervised fine-tuning (SFT) [81], reinforcement learning with human feedback (RLHF) [63, 19], and direct preference optimization (DPO) [66]. Aligned models [82, 25, 92] produce safe, helpful responses and refuse harmful queries.
@@
Recent works study the vulnerability of aligned language models to jailbreaking attacks, which induce models to generate harmful responses. As these models are accessible to any user prompt, most jailbreaks are prompt-based [101, 53, 27, 95], crafting inputs that bypass alignment by compelling models to respond to queries they would otherwise refuse. Recently, some works also present parameter-based jailbreaks where an attacker tunes billions of parameters of a model on a small set of harmful text to remove alignment without altering the input [93, 4]. In contrast, our attack exploits Rowhammer [42] to flip at most 25 bits in a model to jailbreak it at runtime.
@@
Half-precision floating-point numbers. Modern computer systems use IEEE754 floating-point format to represent model parameters. Unlike conventional models that typically use single-precision (32-bit), language models with billions of parameters are commonly deployed in half-precision (16-bit), with a sign bit, 5 exponent bits, and 10 mantissa bits to reduce memory footprint. For instance, in half-precision, is represented as , where is the sign, 1.75 is the mantissa, and is the exponent. Here, the exponent bits play a significant role: flipping a bit in the mantissa induces a marginal change to , whereas a bit-flip in the exponent can shift the value to . Our attacker favors this insight to jailbreak a model with a few bit-flips.
@@
Rowhammer is a software-based fault injection attack [42]. It exploits DRAM disturbance errors by repeatedly accessing DRAM rows (i.e., aggressors) to accelerate charge leakage in adjacent rows, leading to a bit-flip. Recent works have shown numerous Rowhammer exploits on both CPUs [46, 75, 37, 26, 36, 43, 31, 61] and GPUs [51]. Rowhammer exploits have also targeted neural networks to achieve performance degradation [67, 94], targeted misclassification [70, 7, 11], backdooring [69, 99, 45, 14, 80, 50], and model extraction [68]. Our work advances this research direction by presenting the first work on jailbreaking language models through Rowhammer.
@@ 3 Threat Model
Setting. We consider an adversary aiming to jailbreak LLMs by tampering with model parameters in memory at runtime. We assume a setting where a victim deploys a user-facing language model in a GPU-sharing environment, such as a machine-learning-as-a-service (MLaaS) platform. These platforms are increasingly common, as few entities have the resources to operate large-scale models and provide LLM services at scale. We assume the model has been safety-aligned and trained to deny harmful queries (e.g., answering “I cannot answer that” to “how to create an explosive product?”).
We assume co-location of workloads from multiple tenants in a time-sliced GPU environment, consistent with prior work [51]. Here, CUDA kernels from different tenants are executed in a time-multiplexed manner while memory allocations persist across slices. This deployment model is common and supported by NVIDIA infrastructure [59, 60]. We adopt a framework similar to Run:AI [58] and assume that GPU memory swapping [23] is enabled: oversubscribed workloads offload idle job states to CPU memory and restore them upon resumption, improving utilization [23] and enabling multiple tenants to run workloads that would otherwise exceed device capacity. Like prior work [51], we assume GPU memory is managed with RAPIDS Memory Manager (RMM) [1], a widely used allocator for CUDA that improves performance compared to the default cudaMalloc[30]. We assume ECC is disabled, which is often the case in practice, since enabling it has up to 6% memory overhead and 10% slowdown [51, 76].
@@
Knowledge. We consider a white-box attacker knowing the target model’s architecture and parameters, as in most prior works [94, 69, 80, 11, 99, 45]. This is realistic, since many services rely on open-source models [57]. But, this knowledge is not strictly required: we show that bit-flips transfer across models (§6.4), enabling black-box exploitation. The attacker requires no training data. The attacker may search for model parameters with knowledge of vulnerable memory locations (hardware-aware) or search for target bits without any hardware constraints (hardware-agnostic). Since prior attacks assume at most one of these scenarios, our assumption enables a more comprehensive study.
@@
Capabilities. Our attacker exploits Rowhammer [42] attacks to induce targeted bit-flips in memory. The adversary co-locates a process alongside the victim’s process on the same time-sliced GPU. The adversary can execute CUDA kernels natively on the GPU with user-level privileges. The host OS is uncompromised, and confinement policies (e.g., process isolation) are enforced. Our attacker can reverse-engineer GPU DRAM address mappings and place target weights at desired physical locations using techniques proposed in prior work [51]. We provide further details on the Rowhammer exploitation process in §7. While we focus on LLMs running on GPUs, our attack can also generalize to CPU-based deployments, similar to prior Rowhammer ML exploits on CPUs [94, 68, 80, 45, 50]; our attack is easier to mount on CPU memories (e.g., DDR4, LPDDR4), which are orders of magnitude more vulnerable than the GDDR6 we consider [51, 36].
@@ 4 PrisonBreak
Here, we introduce PrisonBreak, our attack that jailbreaks LLMs with targeted bit-flips. Figure 1 illustrates its workflow:
Offline procedure. The attacker identifies target bit locations in the victim model’s memory using a progressive bit-search. First, the attacker analyzes the victim model and identifies critical weights whose perturbations are most effective in jailbreaking. Next, each bit in the floating-point representation of a critical weight is flipped one at a time in isolation to identify the most effective bit-flip. This bit is selected and flipped in the model, and it persists for the rest of the search. The attacker repeats these steps till it reaches an attack success rate of, e.g., 80–95%, as in prior work [101].
PrisonBreak supports two modes: hardware-agnostic and hardware-aware search, depending on the attacker’s knowledge (see §3). In the hardware-agnostic mode, the attacker templates memory only after running PrisonBreak. This approach is viable on highly vulnerable DRAM (e.g., LPDDR4 [36]), and represents an upper bound on attack success. We test this mode in §6. In the hardware-aware mode, the attacker first profiles memory, and PrisonBreak can only use bits identified as vulnerable. This is required for DRAM that is more resilient to Rowhammer (e.g., some DDR4 [36] modules and GDDR6 on GPUs [51]). Using this mode, we demonstrate successful jailbreaks using only two vulnerable bit locations in §7. We detail both modes in §5.3.
Online procedure. After identifying target bits, PrisonBreak jailbreaks the victim language model by flipping them. Following our threat model in §3, this exploitation is accomplished using advanced Rowhammer techniques, including memory templating [64] and massaging [46], and advanced Rowhammer attacks [26, 36, 37] that bypass in-DRAM target row refresh and induce precise bit-flips.